OpenStudy (anonymous):

complex fourth root of -16

ganeshie8 (ganeshie8):

As a start : express -16 in trig/polar form

ganeshie8 (ganeshie8):

can you try?

OpenStudy (anonymous):

4\[4\sqrt{-16[\cos(\theta/4+2k \pi/4)+isin(\theta/4+2k \pi/4)}\]

ganeshie8 (ganeshie8):

looks latex has some problem

ganeshie8 (ganeshie8):

\[-16 = -16+0i\]